OpenStudy (anonymous):

In a typing class, the average number of words per minute N typed after t weeks of lessons was found to be N= 154/(1+ 6.4e^-.2t). Find the time necessary to type 65 words per minute.

OpenStudy (anonymous):

\[65=\frac{154}{1+ 6.4e^{-.2t}}\] is a start

OpenStudy (anonymous):

ok thanks but what do i do next

OpenStudy (anonymous):

i would start by writing \[1+6.4e^{-.2t}=\frac{154}{65}\]

OpenStudy (anonymous):

wait so you multiplied 1+6.4e^-.2e and then divided by 65?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\(a=\frac{b}{c}\iff c=\frac{b}{a}\)

OpenStudy (anonymous):

then subtract 1 from both sides

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

you should get \[e^{-.2t}=\frac{89}{65}\] if my arithmetic is right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait what happened to the 6.4 did you divide?

OpenStudy (anonymous):

i still don't fully understand but thank you for your help

OpenStudy (anonymous):

oh damn hold on

OpenStudy (anonymous):

\[e^{-.2t}=\frac{89}{65}\] is a mistake sorry

OpenStudy (anonymous):

should be \(6.4e^{-2t}=\frac{89}{65}\)

OpenStudy (anonymous):

so you next have to divide by \(6.4\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

numbers get real ugly, time to revert to decimals

OpenStudy (anonymous):

maybe \[e^{-.2t}=.235577\]

OpenStudy (anonymous):

then write in logarithmic form as\[-.2t=\ln(.23557)\] and finally divide by \(-.2\) use a calculator

OpenStudy (anonymous):

check my arithmetic, it is not perfect

OpenStudy (anonymous):

ok and thanks again :)